OpenStudy (anonymous):

solving systems by substitution: -7x - 2y = -13 x - 2y=11 Help! (more questions)

hero (hero):

Multiply the second equation by -1 to get -x + 2y = -11 Now add the equations together to get -8x = -24 Next, swap the sides to get: 8x = 24 Now divide both sides by 8: x = 24/8 x = 3 Solve for y on your own.